41 FIND ALL VALVES OF Of [0,2T] WHERE THE CURVE r = 1-SINO HAS A HORIZONTAL TANGENT LINE, 8 667. 162
Answer #1

solution given that raksino The tangent line is given by dy daldo flossimoff of foro flos coso - Helsino 17 - daro for HorizTI - 2 sino cos of coso = 6 Coso (1-2 sno)=o o ņ colo to 1- 2 sino = 0 Simo: a/ O E [ 01 29 Loso=0 o T/2, 35/2 and sin o= 2 =